# Swarming Lattice in the Frustrated Vicsek-Kuramoto Model

### v2

@@ Abstract
We introduce a frustration parameter $\alpha$ into the Vicsek-Kuramoto model of self-propelled particles. While the system exhibits conventional synchronized states, such as global phase synchronization and swarming, for low frustration ($\alpha<\pi/2$), beyond the critical point $\alpha=\pi/2$, a Hopf–Turing bifurcation drives a transition to a resting hexagonal lattice, accompanied by spatiotemporal patterns such as vortex lattices and dual-cluster lattices with oscillatory unit-cell motions. Lattice dominance is governed by coupling strength and interaction radius, with a clear parametric boundary balancing pattern periodicity and particle dynamics. Our results demonstrate that purely orientational interactions are sufficient to form symmetric lattices, challenging the necessity of spatial forces and illuminating the mechanisms driving lattice formation in active matter systems.
@@
Since the 2010s, the Vicsek model has been linked to the Kuramoto model [RevModPhys.77.137, RODRIGUES20161, 2019Emergence] by equating motion direction with oscillator phase [PhysRevLett.108.248101, doi:10.1142/S0218202513400095], forming a Vicsek-Kuramoto framework. Recent models combining self-propulsion and phase interactions reveal chirality [PhysRevLett.119.058002, PhysRevLett.127.238001, LU2025115794, Lu_2025], memory effects [Sumino2012, PhysRevLett.114.168001], chimeras [PhysRevE.98.032219, PhysRevE.102.022604], and nematic [Sumino2012, PhysRevLett.114.168001, PhysRevLett.127.238001] or anti-aligning interactions [PhysRevE.109.024602], often self-organizing into complex patterns. Evidence suggests purely orientational interactions can form symmetric structures; for example, microtubules form vortex lattices via nematic alignment [Sumino2012], modeled with angular memory [PhysRevLett.114.168001], and anti-aligning interactions can yield traveling hexagonal lattices [PhysRevE.109.024602]. However, these often require specific interactions or effects.
@@
for $i=1,2,\ldots,N$, where $N$ is the population size, $v$ is self-propulsion speed, $\mathbf{p}\left(\theta_{i}\right)=\left(\cos\theta_{i},\sin\theta_{i}\right)$ is the unit vector in the direction $\theta_{i}$, $\omega_{i}$ is the intrinsic frequency, $A_{i}\left(t\right)=\left\{j:\left|\mathbf{r}_{i}\left(t\right)-\mathbf{r}_{j}\left(t\right)\right|\leqslant d_{0}\right\}$ is the set of neighbors within coupling radius $d_{0}$ and $|A_{i}|$ its cardinality, $K\left(\geqslant 0\right)$ is coupling strength, $\bm{\xi}$ and $\eta$ are Gaussian white noises with zero mean and unit variance, and $D$ and $D_{r}$ are translational and rotational diffusion coefficients, respectively. Here, the interaction function is defined as $F\left(\theta\right)=\sin\left(\theta+\alpha\right)-\sin\alpha$, where $\alpha$ is frustration. The term $\sin\alpha$ ensures synchronization ($\theta_{j}=\theta_{i}$) remains an equilibrium [10.1143/PTP.79.1069]. This model generalizes alignments [PhysRevLett.119.058002, PhysRevResearch.1.023026, Escaff2020, PhysRevLett.127.238001, PhysRevLett.133.258302], anti-alignments [PhysRevE.109.024602, PhysRevE.110.024603], and self-propelled chimera [PhysRevE.98.032219, PhysRevE.102.022604]. It reduces to normal Vicsek–Kuramoto for $\alpha=0$ and anti-aligning interactions for $\alpha=\pi$. We set $\omega_{i}=0$ (identical particles) for simplifying the theoretical analysis and $D=D_{r}=0$ (deterministic dynamics) for clarity; simulations with heterogeneous frequencies and noise are provided in the Supplemental Material [mycomment2023], which manifest qualitatively similar behaviors.
@@
Conclusions. Frustrated alignment in the Vicsek-Kuramoto model can induce a resting hexagonal lattice structure in self-propelled particles, demonstrating that purely orientational interactions are sufficient for symmetric pattern formation without explicit spatial forces. This system exhibits a rich array of collective states: For low frustration parameters ($\alpha<\pi/2$), it achieves synchronized states such as global synchronization and swarming, whereas beyond the critical point ($\alpha>\pi/2$), it transitions into lattice patterns including vortex lattices and dual-cluster lattices, as visualized in the spatial configurations (Fig. 1). The critical transition at $\alpha=\pi/2$ is characterized by a Hopf-Turing bifurcation, leading to spatiotemporal patterns with respiration-like dynamics within the unit cells (Fig. 2). The dominance of the lattice structure depends on the coupling strength and radius, with a clear boundary in the $(K,d_{0})$ parameter space (Fig. 3), explained by balancing pattern periodicity and particle dynamics (Eq. (15)). This work establishes frustration as a design principle for tuning lattice topology in active matter systems, but leaves open questions regarding the stability of pattern selection and the prediction of respiration frequencies, warranting further investigation to characterize the nature of these emergent structures.

### v3

# Swarming Lattice in Frustrated Vicsek-Kuramoto Systems

@@ Abstract
We introduce a frustration parameter $\alpha$ into the Vicsek-Kuramoto systems of self-propelled particles. While the system exhibits conventional synchronized states, such as global phase synchronization and swarming, for low frustration ($\alpha<\pi/2$), beyond the critical point $\alpha=\pi/2$, a Hopf–Turing bifurcation drives a transition to a resting hexagonal lattice, accompanied by spatiotemporal patterns such as vortex lattices and dual-cluster lattices with oscillatory unit-cell motions. Lattice dominance is governed by coupling strength and interaction radius, with a clear parametric boundary balancing pattern periodicity and particle dynamics. Our results demonstrate that purely orientational interactions are sufficient to form symmetric lattices, challenging the necessity of spatial forces and illuminating the mechanisms driving lattice formation in active matter systems.
@@
Since the 2010s, the Vicsek model has been linked to the Kuramoto model [RevModPhys.77.137, RODRIGUES20161, 2019Emergence] by equating motion direction with oscillator phase [PhysRevLett.108.248101, doi:10.1142/S0218202513400095], forming a Vicsek-Kuramoto framework. Recent systems combining self-propulsion and phase interactions reveal chirality [PhysRevLett.119.058002, PhysRevLett.127.238001, LU2025115794, Lu_2025], memory effects [Sumino2012, PhysRevLett.114.168001], chimeras [PhysRevE.98.032219, PhysRevE.102.022604], and nematic [Sumino2012, PhysRevLett.114.168001, PhysRevLett.127.238001] or anti-aligning interactions [PhysRevE.109.024602], often self-organizing into complex patterns. Evidence suggests purely orientational interactions can form symmetric structures; for example, microtubules form vortex lattices via nematic alignment [Sumino2012], modeled with angular memory [PhysRevLett.114.168001], and anti-aligning interactions can yield traveling hexagonal lattices [PhysRevE.109.024602]. However, these often require specific interactions or effects.
@@
for $i=1,2,\ldots,N$, where $N$ is the population size, $v$ is self-propulsion speed, $\mathbf{p}\left(\theta_{i}\right)=\left(\cos\theta_{i},\sin\theta_{i}\right)$ is the unit vector in the direction $\theta_{i}$, $\omega_{i}$ is the intrinsic frequency, $A_{i}\left(t\right)=\left\{j:\left|\mathbf{r}_{i}\left(t\right)-\mathbf{r}_{j}\left(t\right)\right|\leqslant d_{0}\right\}$ is the set of neighbors within coupling radius $d_{0}$ and $|A_{i}|$ its cardinality, $K\left(\geqslant 0\right)$ is coupling strength, $\bm{\xi}$ and $\eta$ are Gaussian white noises with zero mean and unit variance, and $D$ and $D_{r}$ are translational and rotational diffusion coefficients, respectively. Here, the interaction function is defined as $F\left(\theta\right)=\sin\left(\theta+\alpha\right)-\sin\alpha$, where $\alpha$ is frustration. The term $\sin\alpha$ ensures synchronization ($\theta_{j}=\theta_{i}$) remains an equilibrium [10.1143/PTP.79.1069]. This system generalizes alignments [PhysRevLett.119.058002, PhysRevResearch.1.023026, Escaff2020, PhysRevLett.127.238001, PhysRevLett.133.258302], anti-alignments [PhysRevE.109.024602, PhysRevE.110.024603], and self-propelled chimera [PhysRevE.98.032219, PhysRevE.102.022604]. It reduces to normal Vicsek–Kuramoto for $\alpha=0$ and anti-aligning interactions for $\alpha=\pi$. We set $\omega_{i}=0$ (identical particles) for simplifying the theoretical analysis and $D=D_{r}=0$ (deterministic dynamics) for clarity; simulations with heterogeneous frequencies and noise are provided in the Supplemental Material [mycomment2023], which manifest qualitatively similar behaviors.
@@
Conclusions. Frustrated alignment in the Vicsek-Kuramoto system can induce a resting hexagonal lattice structure in self-propelled particles, demonstrating that purely orientational interactions are sufficient for symmetric pattern formation without explicit spatial forces. This system exhibits a rich array of collective states: For low frustration parameters ($\alpha<\pi/2$), it achieves synchronized states such as global synchronization and swarming, whereas beyond the critical point ($\alpha>\pi/2$), it transitions into lattice patterns including vortex lattices and dual-cluster lattices, as visualized in the spatial configurations (Fig. 1). The critical transition at $\alpha=\pi/2$ is characterized by a Hopf-Turing bifurcation, leading to spatiotemporal patterns with respiration-like dynamics within the unit cells (Fig. 2). The dominance of the lattice structure depends on the coupling strength and radius, with a clear boundary in the $(K,d_{0})$ parameter space (Fig. 3), explained by balancing pattern periodicity and particle dynamics (Eq. (15)). This work establishes frustration as a design principle for tuning lattice topology in active matter systems, but leaves open questions regarding the stability of pattern selection and the prediction of respiration frequencies, warranting further investigation to characterize the nature of these emergent structures.